OpenStudy (anonymous):

2 7 3 8 4 9 5 __ what number comes nxt?

OpenStudy (anonymous):

6

OpenStudy (igreen):

Add 5, subtract 4.

OpenStudy (igreen):

Answer is 10

OpenStudy (anonymous):

Your answer is 6

OpenStudy (igreen):

2 + 5 = 7 7 - 4 = 3 3 + 5 = 8 8 - 4 = 4 4 + 5 = 9 9 - 4 = 5 5 + 5 = 10

OpenStudy (anonymous):

lml thxs @iGreen

OpenStudy (igreen):

Np

OpenStudy (igreen):

\(\bbox[5pt, #000000 ,border:5px solid #00FF00]{\bf\huge \color{#00FF00}\sim \! \! \! \! \color{#00FF00}i\it\Huge\color{#00FF00}G\color{#00FF00}r\color{#00FF00}e\color{#00FF00}e\color{#00FF00}n\bf\color{#00FF00}ツ}\)

OpenStudy (igreen):

There it goes..lol.